Three equal masses P,Q and R are pulled with a constant force F. They are connected to each other with strings. the ratio of the tension between PQ and QR is?

Grade:12

3 Answers

rah
7 years ago
pls give the direction of applied forces. if the forces are at equal angles of 120 and on a leveled surface, then the force will be f= masin120...
Riya
3 years ago
Take
P=Q=R=m
 Tension btw Pand Q as T1
Tension  btw Q and R as T2
   For p,
 F-T1=ma.....equ 1
    For Q
T1-T2=ma.......equ 2
    For R
T2=ma.....equ 3
Sub equ 3 in equ 2 to find tension btw PQ(T1)
T1-ma=ma
T1=2ma
So
T1:T2
PQ:QR
2:1
 
     
Yashraj
3 years ago
  1. Practice previous year questions many times..
  2. Appear for a test series and analyse it properly (Giving 2–3 hours)
  3. Revise your notes multiple times till they enter your soul :)
  4. Find your strong points and chapters, work on them..
  5. Solve other people’s doubts. It makes your own concepts very clear